# Affine Chabauty I

### v1

@@ 6.4. Effective bounds on the number of $S$-integral points
The idea is to bound $\mathcal{Y}(\mathbb{Z}_{S})_{\Sigma}$ via (6.1) for each $S$-integral reduction type and sum over all $\Sigma$. Let us first consider the situation with fixed prime-to-$p$ reduction type $\Sigma^{\prime}\in\prod_{\ell\neq p}\mathcal{C}_{\ell}$. Let $\mathcal{Y}(\mathbb{Z}_{S})_{\Sigma^{\prime}}$ be the set of $S$-integral points whose prime-to-$p$ reduction type is equal to $\Sigma^{\prime}$. For a multiplicity-1 component $C\subseteq\mathcal{X}_{p}$ with $(C^{\mathrm{sm}}\cap\mathcal{Y})(\operatorname{\mathbb{F}}_{p})\neq\emptyset$, write $\Sigma=(\Sigma^{\prime},C)$ for the reduction type whose prime-to-$p$ part is $\Sigma^{\prime}$ and whose $p$-part is $\Sigma_{p}=C^{\mathrm{sm}}\smallsetminus\operatorname{\mathcal{D}}$. Since $p\not\in S$, the cuspidal part of $\Sigma$ (see §5.2) is determined by $\Sigma^{\prime}$. The Affine Chabauty Condition (1.1) implies (5.4), so by ˜5.3 there exist a non-zero log differential $\omega=\omega(\Sigma^{\prime})\in\mathrm{H}^{0}(X_{\mathbb{Q}_{p}},\Omega^{1}(D))$ and constants $c_{\Sigma}\in\mathbb{Q}_{p}$ such that $\int_{P_{0}}^{P}\omega=c_{\Sigma}$ for all $P\in\mathcal{Y}(\mathbb{Z}_{S})_{\Sigma}$ of reduction type $\Sigma=(\Sigma^{\prime},C)$. Let $n_{C}$ be the number of zeros of $\widetilde{\omega}^{C}$ in $(C^{\mathrm{sm}}\cap\mathcal{Y})(\operatorname{\mathbb{F}}_{p})$ counted with multiplicity, so we have
@@
Similar to ˜5.3, the bound in ˜C can be improved by observing that different reduction types can lead to the same combination $(\omega,c)$ and hence to the same Chabauty function $\rho$ as in (1.3). Namely, let $T$ denote the set of all primes $\ell$ for which the horizontal divisor $\operatorname{\mathcal{D}}$ intersects $\mathcal{X}_{\ell}$ in more than one component. Note that $T$ is a subset of the set of primes of bad reduction for $\mathcal{X}$. In particular, $T$ is finite. We can assume that $(\mathcal{X}_{\ell}^{\mathrm{sm}}\cap\operatorname{\mathcal{D}})(\operatorname{\mathbb{F}}_{\ell})\neq\emptyset$ for all $\ell\in S$. If two $S$-integral reduction types $\Sigma$ and $\Sigma^{\prime}$ have the same cuspidal part $\Sigma^{c}=(\Sigma^{c}_{\ell})_{\ell\in S_{0}}$ (with $S_{0}\subseteq S$) and satisfy $\Sigma_{\ell}=\Sigma^{\prime}_{\ell}$ for every $\ell\in T$, then they give rise to the same pair $(\omega,c)$. Indeed, by ˜2.16 we then have $\mathfrak{S}_{\ell}(P_{0},\Sigma_{\ell})=\mathfrak{S}_{\ell}(P_{0},\Sigma^{\prime}_{\ell})$ for all primes $\ell$, so the same $(\omega,c)$ by ˜5.2. Moreover, by ˜2.31, the component reduction types at primes in $S\smallsetminus T$ are already covered by cuspidal reduction types, so it is enough to consider sets $S_{0}\subseteq S$ containing $S\smallsetminus T$, while at primes $\ell$ in $S\cap T$, only those components not containing a point in $(\mathcal{X}_{\ell}^{\mathrm{sm}}\cap\operatorname{\mathcal{D}})(\operatorname{\mathbb{F}}_{\ell})$ need to be considered separately. This shows that the bound of ˜C can be improved to
@@
where we set $n_{\ell}^{\prime}=n_{\ell}$ for $\ell\not\in S$ but define $n_{\ell}^{\prime}$ as the number of components of $\mathcal{X}_{\ell}$ containing at least one smooth $\operatorname{\mathbb{F}}_{\ell}$-point but no smooth $\operatorname{\mathbb{F}}_{\ell}$-point on $\operatorname{\mathcal{D}}$ for $\ell\in S$. In the case $S=\emptyset$ the formula simplifies to
$\mathcal{Y}(\mathbb{Z})\leq(\#\mathcal{Y}^{\mathrm{sm}}_{p}(\operatorname{\mathbb{F}}_{p})+2g-2+n)\prod_{\ell\in T\smallsetminus\{p\}}n_{\ell}.$ (6.4)
@@
Our results can be used to bound the number of solutions to this equation in $\mathbb{Z}[1/q]$ for any fixed prime $q\neq 3$. A smooth compactification of $Y=\mathcal{Y}_{\mathbb{Q}}$ is the rank 1 elliptic curve $X:Y^{\prime 2}+7Y^{\prime}Z^{\prime 3}=X^{\prime 3}-Z^{\prime 3}$, into which $Y$ embeds via $(x,y)\mapsto[y:1:x]$. Let $\mathcal{X}/\mathbb{Z}$ be the Weierstrass model defined by the same equation. This model is regular, the primes of bad reduction are $3$ and $5$, and both bad fibres have only one component. In the coordinates $x^{\prime}=X^{\prime}/Z^{\prime}$ and $y^{\prime}=Y^{\prime}/Z^{\prime}$, the divisor of cusps in $X$ is $D=\{y^{\prime}=0\}$, consisting of the rational point $Q_{1}=(1,0)$ and the quadratic point $Q_{2}$ with $Q_{2}({\overline{\mathbb{Q}}})=\{(\zeta_{3},0),(\zeta_{3}^{-1},0)\}$, where $\zeta_{3}$ is a primitive third root of unity. The horizontal closures $\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$ are normal and only intersect in the fibre over $3$, so the model $\mathcal{X}$ is $D$-transversal at all primes except $3$. We take $S=\{q\}$ for any prime $q\neq 3$. Let $\Sigma$ be an $S$-integral reduction type. By ˜2.31 it is enough to consider reduction types with $\Sigma_{q}$ cuspidal. The number of such reduction types is $\operatorname{\mathcal{D}}(\operatorname{\mathbb{F}}_{q})$, which is 1 or 3, depending on whether $q$ splits in $\mathbb{Q}(\zeta_{3})$. The Affine Chabauty Condition (1.2) is satisfied:
@@
Thus, by ˜6.6, there are at most $\mathcal{Y}(\operatorname{\mathbb{F}}_{p})+3$ $S$-integral points of reduction type $\Sigma$, for any prime $p>2g+n=5$. Taking $p=7$, we have $\mathcal{Y}(\operatorname{\mathbb{F}}_{p})=3$, so we obtain the bound $\mathcal{Y}(\mathbb{Z}[1/q])_{\Sigma}\leq 6$. Multiplying with the number of reduction types which are cuspidal at $q$, we obtain
$\mathcal{Y}(\mathbb{Z}[\tfrac{1}{q}])\leq\begin{cases}6,&\text{if $q\equiv-1\bmod 3$},\\ 18&\text{if $q\equiv+1\bmod 3$}.\end{cases}$
@@ 7. Application: even degree hyperelliptic curves
7.1. Hyperelliptic curves over $\mathbb{Q}$
Let $\mathcal{Y}/\mathbb{Z}$ be a hyperelliptic curve defined by an affine equation $y^{2}=f(x)$ with $f\in\mathbb{Z}[x]$ squarefree of degree $2g+2$ such that the leading coefficient of $f$ is a square in $\mathbb{Z}$. Suppose that the Mordell–Weil rank $r$ equals $g$. Let $p$ be a prime of good reduction. In [GM25], the authors use $p$-adic heights to construct a nontrivial function $\rho\colon\mathcal{Y}(\mathbb{Z}_{p})\to\mathbb{Q}_{p}$ and a finite set $T$ such that $\rho(\mathcal{Y}(\mathbb{Z}))\subseteq T$. Such functions can also be obtained with our methods, without the need to compute $p$-adic heights above $p$ for some choice of complementary subspace of the holomorphic differentials in $\mathrm{H}^{1}_{\operatorname{dR}}(X/\mathbb{Q}_{p})$. We can also give good bounds in concrete examples, as the following application illustrates.
@@ 7.1. Hyperelliptic curves over $\mathbb{Q}$
Let $f=x^{2g+2}+f_{2g+1}x^{2g+1}+\dots+f_{0}\in\mathbb{Z}[x]$ be squarefree with $2g+2\geq 4$. Assume that the reduction of $f$ modulo every odd prime is not a square. Assume that either the coefficient $f_{2g+1}$ is odd, or that $f=4P+Q^{2}$ with $P,Q\in\mathbb{Z}[x]$ and $Q$ monic of degree $g+1$ such that the reduction of $y^{2}+Q(x)y-P(x)\in\mathbb{Z}[x][y]$ modulo $2$ does not have a zero $y_{0}\in\operatorname{\mathbb{F}}_{2}[x]$ with $\deg(y_{0})\leq g+1$. Let $p>2g+2$ be a prime such that the reduction of $f$ modulo $p$ remains separable. Assume that the hyperelliptic curve given by $y^{2}=f(x)$ over $\mathbb{Q}$ has Mordell–Weil rank equal to $g$. Then the affine equation
@@
has at most $\mathcal{Y}(\operatorname{\mathbb{F}}_{p})+2g$ integral solutions.

### v2

@@ 6.4. Effective bounds on the number of $S$-integral points
The idea is to bound $\mathcal{Y}(\mathbb{Z}_{S})_{\Sigma}$ via (6.1) for each $S$-integral reduction type and sum over all $\Sigma$. Let us first consider the situation with fixed prime-to-$p$ reduction type $\Sigma^{\prime}\in\prod_{\ell\neq p}\mathcal{C}_{\ell}$. Let $\mathcal{Y}(\mathbb{Z}_{S})_{\Sigma^{\prime}}$ be the set of $S$-integral points whose prime-to-$p$ reduction type is equal to $\Sigma^{\prime}$. For a multiplicity-1 component $C\subseteq\mathcal{X}_{p}$ with $(C^{\mathrm{sm}}\cap\mathcal{Y})(\operatorname{\mathbb{F}}_{p})\neq\emptyset$, write $\Sigma=(\Sigma^{\prime},C)$ for the reduction type whose prime-to-$p$ part is $\Sigma^{\prime}$ and whose $p$-part is $\Sigma_{p}=C^{\mathrm{sm}}\smallsetminus\operatorname{\mathcal{D}}$. Since $p\not\in S$, the cuspidal part of $\Sigma$ (˜3.10) is determined by $\Sigma^{\prime}$. The Affine Chabauty Condition (1.1) implies (5.1), so by ˜5.1 there exist a non-zero log differential $\omega=\omega(\Sigma^{\prime})\in\mathrm{H}^{0}(X_{\mathbb{Q}_{p}},\Omega^{1}(D))$ and constants $c_{\Sigma}\in\mathbb{Q}_{p}$ such that $\int_{P_{0}}^{P}\omega=c_{\Sigma}$ for all $P\in\mathcal{Y}(\mathbb{Z}_{S})_{\Sigma}$ of reduction type $\Sigma=(\Sigma^{\prime},C)$. Let $n_{C}$ be the number of zeros of $\widetilde{\omega}^{C}$ in $(C^{\mathrm{sm}}\cap\mathcal{Y})(\operatorname{\mathbb{F}}_{p})$ counted with multiplicity, so we have
@@
The bound in ˜C can be improved by observing that different reduction types can lead to the same combination $(\omega,c)$ and hence to the same Chabauty function $\rho$ as in (1.3). Namely, let $T$ denote the set of all primes $\ell$ for which the horizontal divisor $\operatorname{\mathcal{D}}$ intersects $\mathcal{X}_{\ell}$ in more than one component. Note that $T$ is a subset of the set of primes of bad reduction for $\mathcal{X}$. In particular, $T$ is finite. We can assume that $(\mathcal{X}_{\ell}^{\mathrm{sm}}\cap\operatorname{\mathcal{D}})(\operatorname{\mathbb{F}}_{\ell})\neq\emptyset$ for all $\ell\in S$. If two $S$-integral reduction types $\Sigma$ and $\Sigma^{\prime}$ have the same cuspidal part and satisfy $\Sigma_{\ell}=\Sigma^{\prime}_{\ell}$ for every $\ell\in T$, then they give rise to the same pair $(\omega,c)$. Indeed, by ˜2.16 we then have $\mathfrak{S}_{\ell}(P_{0},\Sigma_{\ell})=\mathfrak{S}_{\ell}(P_{0},\Sigma^{\prime}_{\ell})$ for all primes $\ell$, so we get the same $(\omega,c)$ by ˜5.2. Moreover, by ˜2.31, at primes in $S\smallsetminus T$ any cuspidal reduction type automatically covers all component reduction types, so it is enough to consider $\Sigma$ with $S\smallsetminus T\subseteq C(\Sigma)$. Moreover, by the same remark, at primes $\ell$ in $S\cap T$ only those components not containing a point in $(\mathcal{X}_{\ell}^{\mathrm{sm}}\cap\operatorname{\mathcal{D}})(\operatorname{\mathbb{F}}_{\ell})$ need to be considered separately. This shows that the bound of ˜C can be improved to
@@
where we set $n_{\ell}^{\prime}=n_{\ell}$ for $\ell\not\in S$ but define $n_{\ell}^{\prime}$ as the number of components of $\mathcal{X}_{\ell}$ containing at least one smooth $\operatorname{\mathbb{F}}_{\ell}$-point but no smooth $\operatorname{\mathbb{F}}_{\ell}$-point on $\operatorname{\mathcal{D}}$ for $\ell\in S$. In the case $S=\emptyset$ this simplifies to
$\mathcal{Y}(\mathbb{Z})\leq(\#\mathcal{Y}^{\mathrm{sm}}_{p}(\operatorname{\mathbb{F}}_{p})+2g-2+n)\prod_{\ell\in T\smallsetminus\{p\}}n_{\ell}^{\prime}.$ (6.4)
We will use the refined bound (6.4) in ˜7.1.
@@
Our results can be used to bound the number of solutions to this equation in $\mathbb{Z}[1/q]$ for any fixed prime $q$. A smooth compactification of $Y=\mathcal{Y}_{\mathbb{Q}}$ is the rank 1 elliptic curve $X:Y^{\prime 2}+7Y^{\prime}Z^{\prime 3}=X^{\prime 3}-Z^{\prime 3}$, into which $Y$ embeds via $(x,y)\mapsto[y:1:x]$. Let $\mathcal{X}/\mathbb{Z}$ be the Weierstrass model defined by the same equation. This model is regular, the primes of bad reduction are $3$ and $5$, and both bad fibres have only one component. In the coordinates $x^{\prime}=X^{\prime}/Z^{\prime}$ and $y^{\prime}=Y^{\prime}/Z^{\prime}$, the divisor of cusps in $X$ is $D=\{y^{\prime}=0\}$, consisting of the rational point $Q_{1}=(1,0)$ and the quadratic point $Q_{2}$ with $Q_{2}({\overline{\mathbb{Q}}})=\{(\zeta_{3},0),(\zeta_{3}^{-1},0)\}$, where $\zeta_{3}$ is a primitive third root of unity. The horizontal closures $\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$ are normal and only intersect in the fibre over $3$, so the model $\mathcal{X}$ is $D$-transversal at all primes except $3$. We take $S=\{q\}$ for a prime $q\neq 3$. (In order to treat the case $q=3$ one could blow up at the point in the mod-3 fibre where $\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$ meet and work with the resulting $D$-transversal model.) Let $\Sigma$ be an $S$-integral reduction type. By ˜2.31 it is enough to consider reduction types with $\Sigma_{q}$ cuspidal. The number of such reduction types is $\operatorname{\mathcal{D}}(\operatorname{\mathbb{F}}_{q})$, which is 1 or 3, depending on whether $q$ splits in $\mathbb{Q}(\zeta_{3})$. The Affine Chabauty Condition (1.2) is satisfied:
@@
Thus, by ˜6.6, there are at most $\mathcal{Y}(\operatorname{\mathbb{F}}_{p})+3$ $S$-integral points of reduction type $\Sigma$, for any prime $p>2g+n=5$ with $p\neq q$. Assume that $q\neq 7$, then taking $p=7$, we have $\mathcal{Y}(\operatorname{\mathbb{F}}_{p})=3$, so we obtain the bound $\mathcal{Y}(\mathbb{Z}[1/q])_{\Sigma}\leq 6$. Multiplying with the number of reduction types which are cuspidal at $q$, we obtain
$\mathcal{Y}(\mathbb{Z}[\tfrac{1}{q}])\leq\begin{cases}6,&\text{if $q\equiv-1\bmod 3$},\\ 18,&\text{if $q\equiv+1\bmod 3$}.\end{cases}$
@@ 7. Application: even degree hyperelliptic curves
7.1. Recovering Linear Quadratic Chabauty
Let $\mathcal{Y}/\mathbb{Z}$ be a hyperelliptic curve defined by an affine equation $y^{2}=f(x)$ with $f\in\mathbb{Z}[x]$ squarefree of degree $2g+2$. Assume that the leading coefficient of $f$ is a square in $\mathbb{Z}$, so that $Y=\mathcal{Y}_{\mathbb{Q}}$ has two rational cusps at infinity. Suppose that the Mordell–Weil rank $r$ equals $g$. Let $p$ be a prime of good reduction and fix $P_{0}\in\mathcal{Y}(\mathbb{Z})$ for simplicity. In [GM25], the authors use $p$-adic heights to construct a nontrivial function $\rho\colon\mathcal{Y}(\mathbb{Z}_{p})\to\mathbb{Q}_{p}$ and a finite set $T$ such that $\rho(\mathcal{Y}(\mathbb{Z}))\subseteq T$. On the other hand, our ˜5.1 (applied to a desingularised model, see §3.4) shows the existence of a log differential $0\neq\omega\in\mathrm{H}^{0}(X_{\mathbb{Q}_{p}},\Omega^{1}(D))$ and finitely many constants $c_{\Sigma}$ indexed by integral reduction types such that $\int_{P_{0}}^{P}\omega=c_{\Sigma}$ holds for all $P\in\mathcal{Y}(\mathbb{Z})_{\Sigma}$. It is natural to ask whether we get the same functions, i.e. whether the function $\rho$ of [GM25] is given by the integral of a log differential. This is indeed the case and becomes clear once we recall the construction of $\rho$.
For each prime $\ell$ we have a local Coleman–Gross height pairing $h_{\ell}$ defined on degree zero divisors on $X_{\mathbb{Q}_{\ell}}$ with disjoint support, taking values in $\mathbb{Q}_{p}$ [CG89]. For $\ell=p$ the pairing depends on a choice of subspace $W\subseteq\mathrm{H}^{1}_{\operatorname{dR}}(X/\mathbb{Q}_{p})$ which is complementary to the space of holomorphic differentials and isotropic with respect to the cup product pairing. The choice is used to associate to each degree zero divisor $E$ on $X_{\mathbb{Q}_{p}}$ a uniquely determined meromorphic differential $\omega_{E}$ with at most simple poles and with residue divisor equal to $E$. The local height pairing at $p$ is then defined as
$h_{p}(E,F)\coloneqq\int_{F}\omega_{E}.$
For $\ell\neq p$ the height pairing is defined in terms of intersection theory on a regular model of $X$ over $\mathbb{Z}_{\ell}$ by
$h_{\ell}(E,F)\coloneqq-\log(\ell)\cdot i_{\ell}(\mathcal{E},\Psi_{\ell}(F))$
where $\mathcal{E}$ is the horizontal extension of $E$ and $\Psi_{\ell}(F)=\mathcal{F}+\Phi_{\ell}(F)$ is as in §2.3. The global Coleman–Gross height pairing is then defined for degree zero divisors with disjoint support on $X$ as the sum of the local height pairings over all places:
$h(E,F)\coloneqq\sum_{\ell}h_{\ell}(E,F).$
The global height pairing vanishes when one of the divisors is principal, so it is well-defined on the Jacobian of $X$. This is not true for the local height pairings, although one can show that when $E$ is supported on cusps, the function $F\mapsto h_{\ell}(E,F)$ is well-defined on the generalised Jacobian $J_{Y}$.
On a hyperelliptic curve $y^{2}=f(x)$ with $\deg(f)=2g+2$ as above we have two $\mathbb{Q}$-rational points at infinity. Denote them by $\infty_{+}$ and $\infty_{-}$ and fix the degree zero divisor $E\coloneqq\infty_{+}$-$\infty_{-}$ as first argument of the height pairings. The function $h(E,-)\colon J(\mathbb{Q})\to\mathbb{Q}_{p}$ is a well-defined homomorphism on the Jacobian. One may assume that the closure of $J(\mathbb{Q})$ has finite index in $J(\mathbb{Q}_{p})$, for otherwise one can find a non-zero holomorphic differential whose integral vanishes on all of $X(\mathbb{Q})$ by the classical Chabauty method. When the closure of $J(\mathbb{Q})$ has finite index in $J(\mathbb{Q}_{p})$, every homomorphism $J(\mathbb{Q})\to\mathbb{Q}_{p}$ is given by the integral of a holomorphic differential, in particular there exists $\omega_{\operatorname{hol}}\in\mathrm{H}^{0}(X_{\mathbb{Q}_{p}},\Omega^{1})$ such that
[…]
At $\ell=p$, the local height $h_{p}(\infty_{+}-\infty_{-},-)$ is given by the integral of a logarithmic differential $\omega_{E}$ with residues $1$ and 1-1 at $\infty_{+}$ and $\infty_{-}$, respectively. Subtracting the local height at $p$ from the global height and taking $F\coloneqq P-P_{0}$, we have
$\int_{P_{0}}^{P}\omega_{\operatorname{hol}}-\int_{P_{0}}^{P}\omega_{E}=\sum_{\ell\neq p}h_{\ell}(\infty_{+}-\infty_{-},P-P_{0})\quad\text{for $P\in Y(\mathbb{Q})$}.$
The expression on the left hand side defines a function $\rho\colon\mathcal{Y}(\mathbb{Z}_{p})\to\mathbb{Q}_{p}$, whereas the right hand side takes only finitely many values for $P\in\mathcal{Y}(\mathbb{Z})$, corresponding to the possible reduction types of $P$.
The function $\rho$ obtained in this way is given as the integral of the log differential $\omega\coloneqq\omega_{\operatorname{hol}}-\omega_{E}$ and thus of the same form as the function obtained from our ˜5.1. Apart from working also for primes $p$ of bad reduction, our approach has the advantage that it yields directly the log differential $\omega$ without the need to determine $\omega_{\operatorname{hol}}$ and $\omega_{E}$ individually. Note that the splitting of $\omega$ into $\omega_{\operatorname{hol}}$ and $\omega_{E}$ depends on the choice of complementary subspace $W\subseteq\mathrm{H}^{1}_{\operatorname{dR}}(X/\mathbb{Q}_{p})$ but $\omega$ itself is unique up to scaling, independently of any choices.
The height-based approach can be generalised to more general affine curves admitting a divisor of degree zero supported in the cusps. This is satisfied whenever $D$ contains at least two closed points, and in this case there are #|D|1\#|D|-1 independent such divisors. On the other hand, our Affine Chabauty Condition (1.1) in the case $r=g$ and $S=\emptyset$ reads $0<\#|D|+n_{2}(D)-1$, so our approach yields more independent Chabauty functions in general and also applies when there is just a single cusp with non-totally real residue field, for example for even-degree hyperelliptic curves $y^{2}=f(x)$ where $f$ has negative leading coefficient. One reason for this is that the local height pairing $h_{\ell}$ at a prime $\ell\neq p$ only captures a total intersection number $i_{\ell}(-,-)$ in the mod-$\ell$ fibre of a regular model, whereas our $D$-intersection map captures the intersection numbers $i_{x}(-,-)$ at individual closed points $x$, which by (2.2) is more refined information. It seems likely that the height-based method corresponds to the Chabauty–Kim method for the quotient of the abelianised fundamental group of $Y$ described in [LLM23, Remark 4.3], which is an extension of the Tate module of $J$ by a Tate representation, whereas our method corresponds to the full abelianisation, which is an extension of the Tate module by an Artin–Tate representation. In line with Dogra’s terminology [Dog25], one might call the weaker variant “skimmed linear Chabauty”.
On the other hand, using the methods of this paper we cannot handle integral points on hyperelliptic curves of odd degree with $r=g$ as in [BBM16]. In this case there is only one point at infinity and the proof of Theorem 3.1 of loc. cit. uses the function $x\mapsto h(x-\infty,x-\infty)$ in which both arguments of the height pairing depend on $x$, leading to functions involving not just single but double Coleman integrals. An “Affine Quadratic Chabauty” method should be able to recover the results of [BBM16] as a special case.
7.2. A bound for integral points on hyperelliptic curves over $\mathbb{Q}$
Consider a hyperelliptic curve $\mathcal{Y}/\mathbb{Z}$ defined by an affine equation $y^{2}=f(x)$ with $f\in\mathbb{Z}[x]$ monic and squarefree of degree $2g+2$ and suppose that the Mordell–Weil rank $r$ equals $g$. As discussed in §7.1, we obtain a non-zero log differential whose integral takes only finitely many values on integral points. We show that in many cases it takes just one value. With the results from §6 this leads to a good bound on $\mathcal{Y}(\mathbb{Z})$.
@@ 7.2. A bound for integral points on hyperelliptic curves over $\mathbb{Q}$
Let $f=x^{2g+2}+f_{2g+1}x^{2g+1}+\dots+f_{0}\in\mathbb{Z}[x]$ be squarefree with $2g+2\geq 4$. Assume that the reduction of $f$ modulo every odd prime is not a square. Assume that either the coefficient $f_{2g+1}$ is odd, or that $f=4P+Q^{2}$ with $P,Q\in\mathbb{Z}[x]$ and $Q$ monic of degree $g+1$ such that the reduction of $y^{2}+Q(x)y-P(x)\in\mathbb{Z}[x][y]$ modulo $2$ does not have a zero $y_{0}\in\operatorname{\mathbb{F}}_{2}[x]$ with $\deg(y_{0})\leq g+1$. Let $p>2g+2$ be a prime such that the reduction of $f$ modulo $p$ remains separable. Assume that the hyperelliptic curve given by $y^{2}=f(x)$ over $\mathbb{Q}$ has Mordell–Weil rank equal to $g$. Then the affine curve
@@
has at most $\mathcal{Y}(\operatorname{\mathbb{F}}_{p})+2g$ integral points.